Hi the section im working on is natural logs Im confused on this problem.
Use the properties of logarithms to simplify each question and solve it. Round to the nearest hundredth.
ln(x+2)+ln(x-2)=0

thanks :)

Answer by stanbon(48535)

ln(x+2)+ln(x-2)=0
ln[(x+2)(x-2)] = 0
ln[x^2-4] = 0
x^2-4 = e^0
x^2-4 = 1
x^2 = 5
x = sqrt5 or x = -sqrt(5)
Only x = sqrt5 is an acceptable answer
because -sqrt(5)-2 is negative and ln(negative) does not exist.
======================
Cheers,
Stan H.